20 12897.1 = 3-5 A Find Perimeter Find Area 13.5ft. 29ft. —

Find Area
Find Perimeter
Answer #1

18 Ft 28 ft 13:5 Hi 7 kraft Raduis of the semicircle = 28-13.5 Base of the triangular part 1 85 6375) = 14.5 ft. = 324 -182:2Area of the semicircle = // 1772 블 =ŽBX1405114.5 = 330-39 Ħ Area of quachant circle = 1972 4 = 143.196 ~ 143.2# Area of the r